OpenStudy (anonymous):

Just started learning about derivatives and the chain rule. Chain rule confuses me, so... heres the problem. f(x)=(5x^2-4x)^3 According to the chain rule, f'(x)=(10x-4)*3*(5x^2-4x)^2 but according to my thinking... n=5x^2-4x therefore, f(x)=n^3 and f'(x)=3n^2 so... f'(x)=3(5x^2-4x)^2

OpenStudy (sriram):

yep but while differentiating n it is valid for n but n is actually a quadratic in x which is also changing hence we use chain rule learn limits u may understand derivatives better

OpenStudy (anonymous):

f'(x)=(10x-4)*3*(5x^2-4x)^2 is correct. if n = 5x^2-4x then the derivative of (n)^3 with respect to x (not n) will be 3* (dn/dx) * n^2

OpenStudy (anonymous):

dn/dx = 10x-4, so in the end it's f'(x)=(10x-4)*3*(5x^2-4x)^2

OpenStudy (anonymous):

what does (dn/dx) mean? like I said, I just started learning this stuff.

OpenStudy (sriram):

think of it like this dx=dn*dx/dn

OpenStudy (sriram):

derivative of n wrt x